OpenStudy (anonymous):

use the FOIL method to solve: (3x-4)(3x+4)

Parth (parthkohli):

\(\begin{align} \text{F.O.I.L. = First Outside Inside Last}\\ \hline \\ &(ax+b)(cx+d)\\ \ \\ \text{First:}&(\color{blue}{ax}+b)(\color{blue}{cx}+d)\to \color{blue}{ax}\cdot\color{blue}{cx}=\color{blue}{acx^2}\\ \text{Outside:}&(\color{green}{ax}+b)(cx+\color{green}{d})\to \color{green}{ax}\cdot\color{green}{d}=\color{green}{adx}\\ \text{Inside:}&(ax+\color{red}{b})(\color{red}{cx}+d)\to \color{red}{b}\cdot\color{red}{cx}=\color{red}{bcx}\\ \text{Last:}&(ax+\color{orange}{b})(cx+\color{orange}{d})\to \color{orange}{b}\cdot\color{orange}{d}=\color{orange}{bd}\\ \hline \\ (ax+b)(cx+d)&=\color{blue}{acx^2}+\color{green}{adx}+\color{red}{bcx}+\color{orange}{bd} \end{align}\)

OpenStudy (anonymous):

i just needed it actually solved

Parth (parthkohli):

Eww my LaTeX was hard work

Parth (parthkohli):

jk it's @nbouscal 's work :)

Parth (parthkohli):

\( \color{Black}{\Rightarrow (3x - 4)(3x + 4) }\) \( \color{Black}{\Rightarrow (3x \cdot 3x) + (3x \cdot 4) + (-4 \cdot 3x) + (-4 \cdot 4) }\)

OpenStudy (anonymous):

9x+12x-12x-16?

Parth (parthkohli):

Yeah 9x^2 is the first term

Parth (parthkohli):

And combine like terms

OpenStudy (anonymous):

ok now when i simplify it even more what is it because wouldnt the two like terms (-12 and 12) cancel each other out

Parth (parthkohli):

Yeah

OpenStudy (anonymous):

so itd be 9x^2-16?

Parth (parthkohli):

Yep

OpenStudy (anonymous):

yay correct :) thanks